As a savings plan for college, when their son Bob was born, the Wilburs deposited $10,000 in an account paying 8% compounded annually. How much will the account be worth when Bob is 18 years old?
$_______________

Answer by stanbon(75887)

As a savings plan for college, when their son Bob was born, the Wilburs deposited $10,000 in an account paying 8% compounded annually. How much will the account be worth when Bob is 18 years old?
A(t) = P(1+(r/n))^(nt)
-----
A(18) = 10,000(1+0.08)^18
----
A(18) = 10,000*3.9960
$ $39960.00
------------
Cheers,
Stan H.
-------------